# ColorSwap: A Color and Word Order Dataset for Multimodal Evaluation

### v1

@@ Abstract
This paper introduces the ColorSwap dataset, designed to assess and improve the proficiency of multimodal models in matching objects with their colors. The dataset is comprised of 2,000 unique image-caption pairs, grouped into 1,000 examples. Each example includes a caption-image pair, along with a “color-swapped” pair. We follow the Winoground schema: the two captions in an example have the same words, but the color words have been rearranged to modify different objects. The dataset was created through a novel blend of automated caption and image generation with humans in the loop. We evaluate image-text matching (ITM) and visual language models (VLMs) and find that even the latest ones are still not robust at this task. GPT-4V and LLaVA score 72% and 42% on our main VLM metric, although they may improve with more advanced prompting techniques. On the main ITM metric, contrastive models such as CLIP and SigLIP perform close to chance (at 12% and 30%, respectively), although the non-contrastive BLIP ITM model is stronger (87%). We also find that finetuning on fewer than 2,000 examples yields significant performance gains on this out-of-distribution word-order understanding task. The dataset is here: https://github.com/Top34051/colorswap.
@@ 1 Introduction
Despite the success, recent studies reveal that vision and language models often struggle to comprehend compositional relationships, particularly in differentiating captions with the same words but different word orders. The Winoground dataset Thrush et al. (2022) demonstrates that, given two captions composed of the same words in a different order, the performance of many well-known models in correctly matching these captions to their respective images is close to random chance. It is not well understood whether known multimodal architectures are even capable, in practice, of learning how to perform well at a balanced word-order-understanding task with small-scale finetuning data. Finetuning has been done on some balanced word-order datasets, but only in settings where half of the images are lacking, as far as we are aware Yuksekgonul et al. (2023).
@@ 2 ColorSwap
In this section, we introduce the ColorSwap dataset, comprised of 1,000 unique examples created by four expert annotators with the aid of generative models. Each example includes a caption paired with an image, along with a color-swapped version of the caption and image. We randomly select 700 examples for training and 300 for evaluation.
@@ Results.
Table 2 outlines the performance of the models without finetuning. BLIP and SigLIP exhibit superior performance compared to CLIP and FLAVA, both of which are around the levels of random chance. Matching with an ITM head also improves the image-text matching performance, especially for BLIP. Also, GPT-4V, despite its status as a leading closed-source model, still exhibits genuine errors on the simple task posed by this dataset. A selection of examples and model responses is included in Appendix E.
@@ 4 Conclusion
Acknowledgements
We thank Max Bartolo for suggesting years ago that we should finetune on Winoground - even though Winoground still seems overwhelmingly difficult for models, we were able to show that models can generalize after finetuning on simpler word-order understanding tasks. We appreciate Diyi Yang’s advice on the framing of our work, and additional experiments to run. We thank Aryaman Arora for insightful discussions that ultimately led to the inclusion of the “The key to the shed is blue.” figure.
@@ Appendix D Finetuning Details
Appendix E Qualitative GPT-4V Evaluation

### v2

@@ Abstract
This paper introduces the ColorSwap dataset, designed to assess and improve the proficiency of multimodal models in matching objects with their colors. The dataset is comprised of 2,000 unique image-caption pairs, grouped into 1,000 examples. Each example includes a caption-image pair, along with a “color-swapped” pair. We follow the Winoground schema: the two captions in an example have the same words, but the color words have been rearranged to modify different objects. The dataset was created through a novel blend of automated caption and image generation with humans in the loop. We evaluate image-text matching (ITM) and visual language models (VLMs) and find that even the latest ones are still not robust at this task. GPT-4V and LLaVA score 72% and 42% on our main VLM metric, although they may improve with more advanced prompting techniques. On the main ITM metric, contrastive models such as CLIP and SigLIP perform close to chance (at 12% and 30%, respectively), although the non-contrastive BLIP ITM model is stronger (87%). We also find that finetuning on fewer than 2,000 examples yields significant performance gains on this out-of-distribution word-order understanding task. The dataset is here: https://github.com/Top34051/colorswap and here: https://huggingface.co/datasets/stanfordnlp/colorswap.
@@ 1 Introduction
Despite the success, recent work reveals that vision and language models often struggle to comprehend fine grained distinctions in images Krojer et al. (2022) and compositional relationships, particularly in differentiating captions with the same words but different word orders Thrush et al. (2022). The Winoground dataset from Thrush et al. demonstrates that, given two captions composed of the same words in a different order, the performance of well-known models in correctly matching these captions to their respective images is close to random chance. It is not well understood whether known multimodal architectures are even capable, in practice, of learning how to perform well at a balanced word-order-understanding task with small-scale finetuning data. Finetuning has been done on some balanced word-order datasets, but only in settings where half of the images are lacking, as far as we are aware Yuksekgonul et al. (2023).
@@ 2 ColorSwap
In this section, we introduce the ColorSwap dataset, comprised of 1,000 unique examples created by four expert annotators (each with two or more years or NLP experience and living in the United States) with the aid of generative models. Each example includes a caption paired with an image, along with a color-swapped version of the caption and image. We randomly select 700 examples for training and 300 for evaluation.
@@ Results.
Table 2 outlines the performance of the models without finetuning. BLIP and SigLIP exhibit superior performance compared to CLIP and FLAVA, both of which are around the levels of random chance. Matching with an ITM head also improves the image-text matching performance, especially for BLIP. Also, GPT-4V, despite its status as a leading closed-source model, still exhibits genuine errors on the simple task posed by this dataset.
In Appendix E, we run an experiment to provide evidence that the models’ poor performance is due to the compositional nature of this task and not simply because the images are generated by diffusion models (and so are out of the pretraining distributions). A selection of examples and model responses is included in Appendix F. We report confidence intervals for these results in Appendix G.
@@ 4 Conclusion
5 Limitations
All of the images in ColorSwap come from diffusion models. The poor performance of models on the dataset could come from the fact that the diffusion images are simply out of distribution, not because these models have issues with color-word compositionality - although we provide some evidence against this in Appendix E. Similarly, the captions may not be “in-distribution” for a variety of reasons. There is a risk of misattributing failure reasons. However, if a model scores well, then we can be assured that the model is able to correctly differentiate between color-swapped images in this particular setting. ColorSwap is additionally a static dataset, meaning that models evaluated on it must be vetted for training data contamination.
6 Ethical Considerations
For all diffusion models and datasets that we used, we believe that our use is consistent with their intended use and licenses. However, the diffusion models that we used are trained on a variety of opaque sources from the internet, and may make use of data without creators’ consent.
@@ Appendix D Finetuning Details
Appendix E Is Low Performance Caused by OOD Diffusion Images?
To better understand whether the poor performance on the ColorSwap dataset is due to the images being AI-generated or the models’ inability to distinguish compositional color relationships, we run an additional experiment. Instead of choosing between an image-caption pair and a color-swapped version, we assess whether the models can distinguish between an image-caption pair and another randomly selected image-caption pair from our dataset. If the models’ poor performance is entirely due to their inability to interpret diffusion-generated images, we would expect them to perform poorly on this task as well. We evaluate some of the worst-performing models from our main experiments, along with BLIP, on this task.
Model & Method Text $\uparrow$ Image $\uparrow$ Group $\uparrow$
Random chance 25.00 25.00 16.67
Contrastive matching
CLIP 100.00 99.67 99.67
FLAVA 99.33 100.00 99.33
BLIP 100.00 100.00 100.00
ITM matching
FLAVA 82.67 97.67 81.00
BLIP 100.00 100.00 100.00
Table 6: Performance of models on distinguishing matching ColorSwap image-caption pairs from randomly selected non-matching pairs. Results above chance are bold.
Table 6 suggests that the models can distinguish between matching and non-matching image-text pairs even when the images are AI-generated. This provides some evidence that the poor performance on the ColorSwap dataset is not due to a specific issue with the out-of-distribution nature of AI-generated images.
Appendix F Qualitative GPT-4V Evaluation
@@
Appendix G Confidence Intervals
We provide confidence intervals for the overall model results on ColorSwap in Table 7.
Model & Method Text $\uparrow$ Image $\uparrow$ Group $\uparrow$
Image-text matching models
Random chance 25.00 25.00 16.67
Contrastive matching
CLIP 35.67 [30.33, 41.00] 14.67 [11.33, 19.33] 11.67 [8.67, 16.00]
FLAVA 35.33 [30.33, 41.33] 25.00 [20.33, 30.00] 15.67 [11.67, 20.00]
BLIP 75.67 [70.67, 80.33] 56.00 [50.33, 61.67] 51.00 [45.33, 56.67]
SigLIP 61.67 [56.00, 67.00] 37.00 [31.67, 42.67] 30.33 [25.33, 35.67]
ITM matching
FLAVA 36.33 [31.33, 42.00] 18.67 [15.00, 23.67] 10.33 [7.67, 14.67]
BLIP 94.67 [91.67, 97.00] 89.00 [85.00, 92.00] 87.33 [83.00, 90.67]
Visual Language Models
Random chance 25.00 25.00 6.25
LLaVAR 27.67 [22.67, 33.00] 25.67 [21.00, 31.00] 8.33 [5.67, 12.00]
LLaVA-1.5 69.67 [64.33, 74.67] 54.33 [48.36, 59.67] 42.00 [36.33, 47.33]
GPT-4V 91.33 [87.67, 94.33] 76.33 [71.33, 81.00] 72.00 [66.67, 77.00]
Table 7: Performance of models on ColorSwap with confidence intervals. Results above chance are bold. Note that random chance is different in the ITM versus visual language model (VLM) cases because VLMs output a binary value and ITM models output a float (an effectively continuous value).